Data for Hermann Corporation are shown below: Selling price Variable expenses Contribution margin Per Unit $ 80 52 Percent ofShould the higher-quality components be used? OYes O No
Answer #1

Current net operating income

= Total Contribution margin - Fixed cost

= (4,600*28) - 76,000 = 52,800

Projected net operating income

= (4,600+25%)*(28-5) - 76,000

= 56,250

Increase = 56,250 - 52,800

Net operating income increases by $3450

Yes